An item is regularly priced at 91. It is now priced at a discount of 95%
off the regular price.

Answer by Edwin McCravy(20081)

95% of 91 = (0.95)(91) = 86.45
91 - 86.45 = 4.55
-------------------------------------------------------
An easier way is that if you take off 95% you leave 5%
5% of 91 = (0.05)(91) = 4.55 

Edwin